1.) -22=11-6a (11 minus 6 times what = negative 22)
2.) -40=m-10m (what minus 10 times what = negative 40)
3.) 13t=35 ( 13 times what = 35)

three different questions im sorry but i have even more than five and i have been going through a lot so can u at least answer all my homework questions i spent nearly 7 hours on it and i still have nothing im so desperate rite now help me

Answer by checkley79(3341)

1.) -22=11-6a (11 minus 6 times what = negative 22)
-6a=-22-11
-6a=-33
a=-33/-6
a=33/6 or 5.5 ans.
2.) -40=m-10m (what minus 10 times what = negative 40)
-40=-9m
m=-40/-9
m=4.44 ans.
3.) 13t=35 ( 13 times what = 35)
t=35/13
t=2.69 ans.